# Register Any Point: Scaling 3D Point Cloud Registration by Flow Matching

### v1

@@ Abstract
Point cloud registration aligns multiple unposed point clouds into a common frame, and is a core step for 3D reconstruction and robot localization. In this work, we cast registration as conditional generation: a learned continuous, point-wise velocity field transports noisy points to a registered scene, from which the pose of each view is recovered. Unlike previous methods that conduct correspondence matching to estimate the transformation between a pair of point clouds and then optimize the pairwise transformations to realize multi-view registration, our model directly generates the registered point cloud. With a lightweight local feature extractor and test-time rigidity enforcement, our approach achieves state-of-the-art results on pairwise and multi-view registration benchmarks, particularly with low overlap, and generalizes across scales and sensor modalities. It further supports downstream tasks including relocalization, multi-robot SLAM, and multi-session map merging.
Figure 1: Our method for scalable multi-view point cloud registration. To register multiple unposed point clouds, prior work typically first performs correspondence matching and then optimizes a pose graph (top-left). In contrast, we introduce a single-stage model that directly generates the registered point cloud via flow matching in Euclidean space (top-right), bypassing the need for explicit correspondence matching and pose graph optimization. Our model generalizes across diverse point cloud data from object-centric, indoor, and outdoor scenarios at scan, sub-map, and map levels (bottom).
@@ 1 Introduction
Point cloud registration is a cornerstone in 3D vision, robotics, and photogrammetry with broad applications, from merging multiple partial 3D scans into a consistent 3D model to localizing sensors in an existing 3D map for downstream tasks, including simultaneous localization and mapping (SLAM) [zhang20243d], 3D reconstruction [dong2020jprs-whutls], and robotic manipulation [pomerleau2015ftr]. Yet, obtaining reliable registration in the wild is a hard problem. Real-world data is sparse, noisy, and non-uniform in density; sensors differ in modality and calibration; overlaps between point clouds can be small, and local matches can be ambiguous [huang2021cvpr-predator, sun2025jprs, an2024pointtr, liu2024cvpr-eyoc].
@@
Prevailing approaches for multi-view point cloud registration follow a two-stage pipeline: align all overlapping pairs of scans, then solve a global pose graph to enforce consistency [schneider2012isprs, huang2021ral-bundle]. Pairwise alignment typically relies on matching local feature correspondences with a robust estimator [fischler1981cacm]. While conceptually appealing, this has two limitations: (i) quadratic complexity: cost scales quadratically with the number of scans due to exhaustive correspondence search across all pairs; and (ii) limited global context: the pairwise stage limits capturing global context, hurting performance under low overlap and incomplete observations. Although specialized modules can improve low-overlap pairwise registration [huang2021cvpr-predator, yao2024iccv-parenet] and some works conduct hierarchical registration [dong2018jprs] or edge selection [wang2023cvpr-sghr] to avoid the quadratic cost, these add complexity while remaining tied to iterative pose-graph refinement sensitive to pairwise alignment errors.
@@
Recent 3D vision research departs from this two-stage pipeline by leveraging feed-forward and generative models. In image-based 3D reconstruction, feed-forward approaches [wang2024cvpr-dust3r] encapsulate the entire structure-from-motion process into a single neural network, directly producing globally consistent poses and dense geometry from a set of images. VGGT [wang2025cvpr-vggt] demonstrates that a large transformer can infer all key 3D attributes, including camera poses and depth maps, from one or many views in a single pass. In the point cloud domain, Rectified Point Flow (RPF) [sun2025neurips] pioneered a generative approach to pose estimation by learning a continuous flow field that moves points from random noise to their assembled target positions for multiple object-centric benchmarks. These findings suggest that a single feed-forward model can holistically reason about multiple partial observations and produce a consistent 3D alignment, given sufficient capacity and training data.
@@
Scaling such single-stage models to large-scale, multi-view 3D registration, however, raises another key challenge: the sampling process does not always yield stable, perfectly rigid predictions, especially in cluttered environments where geometry is more diverse than in object-centric settings. Even with an explicit projection step of the final prediction onto $\mathrm{SE}(3)$, as in RPF [sun2025neurips], the post-hoc correction cannot constrain the entire flow trajectory and, thus, the sampled flows can drift away from the flow distribution on which the model was trained, limiting performance.
@@
This motivates our work, a scalable generative model that aligns multiple point clouds in a single stage while explicitly enforcing rigidity. Instead of exhaustive pairwise pose estimation, the model learns to transform all input point clouds directly into a canonical coordinate frame, effectively fusing them into a coherent scene. To make generation robust and satisfy rigid constraints, we propose using rigidity both as a guidance signal for flow sampling and as a criterion for selecting generations. To train at scale, we curate over 100K samples from 17 diverse datasets spanning object-centric, indoor, and outdoor settings. Supervising in Euclidean space across this mixture of data provides strong scene priors that enable the model to complete partial views and generalize across scales and sensor modalities. We will release our code and models at https://github.com/PRBonn/RAP.
In summary, our contributions are four-fold:
  • We propose a generative flow-matching model that performs multi-view point cloud registration in a single stage, bypassing iterative pose graph optimization.
  • We introduce a rigidity-forcing sampling and selection strategy that enforces per-scan rigid constraints and improves registration accuracy.
@@
  • We develop a large-scale training recipe that aggregates over 100K samples from 17 heterogeneous datasets, enabling strong generalization across diverse scenarios, scales, and sensor modalities.
  • We demonstrate robust performance on both pairwise and multi-view benchmarks, achieving substantial improvements over state-of-the-art methods in challenging large-scale, low-overlap scenarios.
@@ 2 Related Work
Pairwise point cloud registration has long relied on local feature matching with robust estimators [fischler1981cacm, yang2020tro, lim2025icra-kissmatcher]. Early approaches rely on hand-crafted local descriptors [rusu2009icra, salti2014cviu] to establish correspondences. Modern methods [bai2020cvpr-d3feat, huang2021cvpr-predator, qin2022cvpr-geotransformer, ao2021cvpr-spinnet, wang2022mm-yoho, liu2024cvpr-eyoc, seo2025iccv, lim2025icra-kissmatcher] replace or augment these with learned local features. Beyond explicit matches, correspondence-free and end-to-end approaches such as RPM-Net [yew2020cvpr] directly supervise the pose with differentiable assignment and iterative refinement. Our approach departs from both families of approaches: rather than seeking correspondences or iteratively refining poses, we learn a conditional velocity field that transports noise to the merged scene, from which rigid transforms are recovered using SVD.
@@
Multi-view point cloud registration is typically handled by first estimating noisy pairwise relative poses and then running pose graph optimization (PGO). In practice, optimization is performed on a factor graph with one node per scan and edges that encode relative pose measurements and their uncertainties under various robust objectives [theiler2015jprs, dong2018jprs, dong2020jprs-whutls, gojcic2020cvpr, wang2023cvpr-sghr, dellaert2012git]. By contrast, our single-stage approach aligns an arbitrary number of scans at once, enforcing multi-view consistency by construction. As a result, our method dispenses with the need for a separate PGO stage and avoids the quadratic costs from pairwise pose estimation. Nonetheless, our predictions can still serve as strong and time-efficient initializations for downstream PGO solvers that incorporate additional signals (e.g., from gravity, IMU, or GNSS) or task-specific constraints.
@@
Generative modeling approaches for 3D data leverage diffusion- and flow-based models to generate geometric structures [nunes2024cvpr, sun2025neurips, sanghi2023cvpr, wang2025iclr-puzzlefusionplusplus, xu2023cvpr-dzts, zeng2024cvpr-ppaw, bian2025iclr, guo2025neurips, zhang2025iccv, du2025cvpr, li2025iccv-garf, ren2024cvpr, nunes2025arxiv, jiang2023neurips-se3diffusion]. Both model generation as stochastic transport from source to target distributions: diffusion via iterative denoising, while flow-matching predicts velocity fields to iteratively transform data. These models have been applied to text-to-shape generation [sanghi2023cvpr, xu2023cvpr-dzts, zeng2024cvpr-ppaw], 3D scene completion [nunes2024cvpr, zhang2025iccv, du2025cvpr], and annotated data generation [ren2024cvpr, nunes2025arxiv, bian2025iclr, guo2025neurips].
More recent works leverage diffusion and flow-matching models to achieve point cloud registration [jiang2023neurips-se3diffusion, sun2025neurips] as a way to overcome limitations of standard approaches. In DiffusionReg [jiang2023neurips-se3diffusion], point cloud registration is formulated as a diffusion process on the SE(3) manifold, generating the corresponding rigid-body transformation between the source and target point clouds. Closer to our approach, RPF [sun2025neurips] uses a conditional flow-matching model to unify pairwise registration and multi-part shape assembly tasks. Our method builds on the point cloud flow matching ideas but differs from RPF in two ways: (i) we do not rely on a specialized overlap prediction network; instead, we represent each point cloud by a sparse set of learned features, enabling scaling to large scenes; and (ii) we rigidify flows at each step to guarantee per-view rigid motion. Together, these choices enable strong generalizations to unseen datasets and robustness at extremely low overlap.
@@ 3 Generative Point Cloud Registration
Figure 2: Overview of our approach to multi-view point cloud registration. Starting with unposed point clouds $\mathcal{P}$, we sample points $\mathcal{Q}$ with corresponding local features $\mathcal{F}$. We use a diffusion transformer with alternating-attention blocks for conditional flow matching that generates the aggregated point cloud $\hat{\mathbf{X}}(0)$ from Gaussian noise $\mathbf{X}(1)$. Finally, we recover the individual transformations $\hat{\mathbf{T}}_{i}$ using SVD from the aggregated point cloud and apply them to the original unposed point clouds to get the registered point clouds $\hat{\mathbf{P}}^{r}$.
@@ 3.1 Problem Definition
We consider the general setting of the multi-view point cloud registration problem. The input is a set of NN ($N\geq 2$) unordered point clouds
$\mathcal{P}:=\left\{\mathbf{P}_{i}\in\mathbb{R}^{3\times M_{i}}\mid i=1,\ldots,N\right\},$ (1)
where $M_{i}$ is the point count of the $i$-th point cloud. These point clouds may come from individual LiDAR or depth-camera scans, or from accumulated point cloud maps built by a SLAM system. We do not assume any initial guess of the transformation from the coordinate frame of each point cloud to a global frame, but we assume that all point clouds are observations of the same scene and can be registered into a single connected point cloud.
The goal is to estimate registered point clouds
$\mathcal{P}^{r}:=\{\mathbf{P}_{i}^{r}\in\mathbb{R}^{3\times M_{i}}\mid i=1,\ldots,N\}$ (2)
in a global coordinate frame. In practice, there may be non-rigid deformations caused by effects such as motion distortion, dynamic objects, or SLAM drift. When these non-rigid effects are negligible, the registered point clouds can be modeled as rigid transformations of the inputs: $\mathcal{T}=\{\mathbf{T}_{i}\in\mathrm{SE}(3)\mid i=1,\ldots,N\}$, which can be recovered via the Kabsch algorithm using SVD between $\mathcal{P}$ and $\mathcal{P}^{r}$.
@@ 3.2 Flow Matching for Multi-view Registration
with $\Delta t=1/\kappa$. After integration, the resulting $\hat{\mathbf{X}}(0)$ approximates the registered point cloud. We then partition $\hat{\mathbf{X}}(0)$ into per-view subsets and estimate the corresponding poses $\hat{\mathbf{T}}_{i}$ via the Kabsch algorithm.
@@ 3.3 Rigidity-forcing Inference
Euler integration with rigidity-forcing.
To illustrate, we define the per-view projection operator $\Pi$ of an estimate $\hat{\mathbf{X}}_{i}(0)$ onto the rigid orbit of the input $\mathbf{P}_{i}$, as
@@
where $(\hat{\mathbf{R}}_{i},\hat{\mathbf{t}}_{i})$ is the optimal rigid transformation between $\mathbf{P}_{i}$ and $\hat{\mathbf{X}}_{i}(0)$, computed via the Kabsch algorithm. Given the current state $\mathbf{X}(t)$ and the velocity $\mathbf{V}_{\theta}(t,\mathbf{X}(t)\!\mid\!\mathbf{C})$, we extrapolate the registered point cloud estimate, as
@@
Rigidity-based sample selection.
In addition to sampling, rigidity also provides a geometric criterion for selecting among multiple generative predictions of the same input. Given a generated estimate $\hat{\mathbf{X}}(0)$ and its rigidified result $\Pi_{{\mathcal{P}}}\!\left(\hat{\mathbf{X}}(0)\right)$, we quantify how well the sample respects per-view rigidity by measuring the average rigidity residual between the original prediction and its rigid projection. For view $i$, let $\hat{\mathbf{X}}_{i}(0)=\{\hat{\mathbf{x}}_{i,j}\}_{j=1}^{M_{i}}$ and $\tilde{\mathbf{X}}_{i}(0)=\Pi_{{\mathbf{P}}_{i}}\!\left(\hat{\mathbf{X}}_{i}(0)\right)=\{\tilde{\mathbf{x}}_{i,j}\}_{j=1}^{M_{i}}$ denote the rigidified points. We define the rigidity residual as
$\mathcal{E}_{\mathrm{rigid}}(\hat{\mathbf{X}}(0)):=\left(\frac{1}{M}\sum_{i}\sum_{j=1}^{M_{i}}\bigl\|\hat{\mathbf{x}}_{i,j}-\tilde{\mathbf{x}}_{i,j}\bigr\|_{2}^{2}\right)^{\!\frac{1}{2}}.$ (8)
At inference, we draw $S$ samples $\{\hat{\mathbf{X}}^{(s)}(0)\}_{s=1}^{S}$ from the flow, compute their corresponding rigidity errors $\mathcal{E}^{(s)}_{\mathrm{rigid}}:=\mathcal{E}_{\mathrm{rigid}}(\hat{\mathbf{X}}^{(s)}(0))$, and select the generation with the lowest rigidity error. This rigidity-based selection is purely geometric and adds negligible overhead to the sampling procedure as it reuses the Kabsch projections already computed during rigidity forcing.

### v2

@@ Abstract
Point cloud registration aligns multiple unposed point clouds into a common reference frame and is a core step for 3D reconstruction and robot localization without initial guess. In this work, we cast registration as conditional generation: a learned, continuous point-wise velocity field transports noisy points to a registered scene, from which the pose of each view is recovered. Unlike prior methods that perform correspondence matching to estimate pairwise transformations and then optimize a pose graph for multi-view registration, our model directly generates the registered point cloud, yielding both efficiency and point-level global consistency. By scaling the training data and conducting test-time rigidity enforcement, our approach achieves state-of-the-art results on existing pairwise registration benchmarks and on our proposed cross-domain multi-view registration benchmark. The superior zero-shot performance on this benchmark shows that our method generalizes across view counts, scene scales, and sensor modalities even with low overlap. [project page]
Figure 1: Our method for multi-view point cloud registration. Prior works perform pairwise registration via correspondence matching and then conduct pose graph optimization (top-left). We use a single-stage flow matching model to generate the registered points (top-middle). Our model generalizes across various view counts, scales, and sensor modalities (bottom) and achieves superior zero-shot performance with the shortest runtime on the cross-domain multi-view registration benchmark (top-right).
@@ 1 Introduction
Point cloud registration is a cornerstone in 3D vision, robotics, and photogrammetry with broad applications, from merging multiple partial 3D scans into a consistent 3D model to localizing sensors in an existing 3D map for downstream tasks, including simultaneous localization and mapping (SLAM) [zhang20243d, pan2021icra-mvls], 3D reconstruction [dong2020jprs-whutls], and robotic manipulation [pomerleau2015ftr]. Yet, obtaining reliable registration in the wild without an initial guess is a hard problem. Real-world data is sparse, noisy, and non-uniform in density; sensors differ in modality and calibration; overlaps between point clouds can be small, and local matches can be ambiguous [huang2021cvpr-predator, sun2025jprs, an2024pointtr, liu2024cvpr-eyoc].
@@
Prevailing approaches for multi-view point cloud registration follow a two-stage pipeline: align all overlapping pairs of scans independently, then solve a global pose graph optimization to enforce consistency [schneider2012isprs, huang2021ral-bundle]. Pairwise alignment typically relies on matching local feature correspondences with a robust estimator [fischler1981cacm, yang2020tro, lim2024ijrr-quatropp, babin2019icra, laserna2025tro]. While conceptually appealing, this has two limitations: (i) quadratic complexity: cost scales quadratically with the number of scans due to exhaustive pairwise registration across all pairs; and (ii) limited global context: the pairwise stage limits capturing global context, hurting performance with low overlap and incomplete observations. Although specialized modules can improve low-overlap pairwise registration [huang2021cvpr-predator, yao2024iccv-parenet] and some works conduct hierarchical registration [dong2018jprs] or edge selection [wang2023cvpr-sghr] to avoid the quadratic cost, these add complexity while remaining tied to iterative pose-graph refinement sensitive to pairwise alignment errors.
@@
Recent 3D vision research departs from this two-stage pipeline by leveraging feed-forward and generative models. In image-based 3D reconstruction, feed-forward approaches [wang2024cvpr-dust3r, leroy2024eccv] encapsulate the entire structure-from-motion process into a single neural network, directly producing globally consistent poses and dense geometry from a set of images. VGGT [wang2025cvpr-vggt] demonstrates that a large transformer can infer all key 3D attributes, including camera poses and depth maps, from one or many views in a single pass. In the point cloud domain, Rectified Point Flow (RPF) [sun2025neurips] pioneered a generative approach to pose estimation by learning a continuous flow field that moves points from random noise to their assembled target positions for multiple object-centric benchmarks. These findings suggest that a single feed-forward model can holistically reason about multiple partial observations and produce a consistent 3D alignment, given sufficient capacity and training data.
@@
Scaling such single-stage models to large-scale, multi-view 3D registration, however, raises another key challenge: the sampling process does not always yield stable, perfectly rigid predictions, especially in cluttered environments where geometry is more diverse than in object-centric settings. Even with an explicit projection step of the final prediction onto $\mathrm{SE}(3)$, as in RPF [sun2025neurips], the post-hoc correction cannot constrain the entire flow trajectory. Thus, the sampled flows can drift away from the flow distribution on which the model was trained, limiting performance.
@@
This motivates our work, a scalable generative model that aligns multiple point clouds in a single stage while explicitly enforcing rigidity. Instead of exhaustive pairwise transformation estimation, the model learns to transform all input point clouds directly into a canonical coordinate frame, effectively fusing them into a coherent scene. To make generation robust and satisfy rigid constraints, we propose using rigidity as a guidance signal for flow sampling at test time. To train at scale, we curate over 100k multi-view registration instances from 17 diverse datasets spanning object-centric, indoor, and outdoor settings. Supervising in Euclidean space across this mixture of data provides strong scene priors that enable the model to complete partial views and generalize across scales and sensor modalities. To address the lack of a unified evaluation protocol for cross-domain generalization, we also introduce a multi-view registration benchmark spanning five scene categories across diverse scales, sensors, and overlap conditions. We will release our code, model, and proposed benchmark to facilitate reproducibility at: https://github.com/PRBonn/RAP.
In summary, our contributions are three-fold:
  • We propose RAP, a generative flow-matching model for single-stage multi-view point cloud registration, together with a rigidity-forcing sampling strategy that enforces per-scan rigid constraints at test time, achieving state-of-the-art performance on both pairwise and multi-view point cloud registration benchmarks.
@@
  • We develop a large-scale training recipe that aggregates over 100k multi-view registration instances from 17 heterogeneous datasets, enabling strong generalization across diverse scenarios, scales, and sensor modalities, including challenging low-overlap conditions.
  • We introduce a challenging cross-domain multi-view registration benchmark spanning five scene categories, on which our method substantially outperforms existing approaches in a zero-shot manner.
@@ 2 Related Work
Pairwise point cloud registration has long relied on local feature matching with robust transformation estimators [fischler1981cacm, yang2020tro, lim2025icra-kissmatcher]. Early approaches rely on hand-crafted local descriptors [rusu2009icra, tombari2010eccv-shot, salti2014cviu, dong2017jprs-bsc, he2016iros, mellado2014cgf-super4pcs] to establish correspondences. Modern methods [choy2019iccv, choy2020cvpr-dgr, gojcic2019cvpr-perfectmatch, bai2020cvpr-d3feat, huang2021cvpr-predator, qin2022cvpr-geotransformer, ao2021cvpr-spinnet, wang2022mm-yoho, wang2023pami-roreg, liu2024cvpr-eyoc, seo2025iccv, yao2024iccv-parenet, wiesmann2022ral-iros, yu2021neurips-cofinet, yu2023cvpr-roitr, li2022cvpr-lepard, shi2021ral, deng2019cvpr, bai2021cvpr-pointdsc, poiesi2023pami-gedi, yu2023cvpr-peal] replace or augment these with learned local features. Beyond explicit matches, correspondence-free and end-to-end approaches [aoki2019cvpr-pointnetlk, wang2019iccv-dcp, yew2020cvpr, wang2019neurips-prnet, bernreiter2021ral-phaser] directly supervise the transformation with differentiable assignment or iterative refinement. Our approach departs from both families: rather than seeking correspondences or iteratively refining poses, we learn a conditional velocity field that transports noise to the merged scene, from which rigid transformations are recovered. Closer to our approach, DeepVCP [lu2019iccv] and DeepPRO [lee2021iccv-deeppro] generate virtual corresponding points near the target, but remain limited to pairwise registration with transformation initial guess or object-scale scenes.
@@
Multi-view point cloud registration is typically handled by first estimating local pairwise relative poses and then synchronizing the transformations via pose graph optimization (PGO) [grisetti2010titsmag, dellaert2012git]. In practice, optimization is performed using a factor graph with one node per scan and edges that encode relative pose measurements and their uncertainties under various robust objectives [theiler2015jprs, choi2015cvpr, dong2020jprs-whutls, gojcic2020cvpr, wang2023cvpr-sghr, jin2024cvpr-multiwaymosaicking, dong2018jprs]. By contrast, our single-stage approach aligns an arbitrary number of scans at once, enforcing multi-view consistency by construction. As a result, our method dispenses with the need for a separate PGO stage and avoids the quadratic costs from pairwise transformation estimation. Nonetheless, our predictions can still serve as strong and time-efficient initializations for downstream PGO solvers that incorporate additional signals (e.g., from gravity, IMU, or GNSS) or task-specific constraints.
@@
Generative modeling approaches for 3D data leverage diffusion- and flow-based models to generate geometric structures [nunes2024cvpr, sun2025neurips, sanghi2023cvpr, wang2025iclr-puzzlefusionplusplus, xu2023cvpr-dzts, zeng2024cvpr-ppaw, bian2025iclr, guo2025neurips, zhang2025iccv, du2025cvpr, li2025iccv-garf, ren2024cvpr, jiang2023neurips-se3diffusion]. Both model generation as stochastic transport from source to target distributions: diffusion via iterative denoising, while flow-matching predicts velocity fields to iteratively transform data. These models have been applied to text-to-shape generation [sanghi2023cvpr, xu2023cvpr-dzts, zeng2024cvpr-ppaw], 3D scene completion [nunes2024cvpr, zhang2025iccv, du2025cvpr], and annotated data generation [ren2024cvpr, nunes2026pami, bian2025iclr, guo2025neurips].
More recent works leverage diffusion and flow-matching models to achieve point cloud registration [jiang2023neurips-se3diffusion, an2024heritagescience, sun2025neurips] as a way to overcome limitations of standard approaches. In DiffusionReg [jiang2023neurips-se3diffusion], point cloud registration is formulated as a diffusion process on the SE(3) manifold, generating the corresponding rigid-body transformation between the source and target point clouds. Closest to our approach, RPF [sun2025neurips] applies Euclidean-space conditional flow matching to pairwise registration and multi-part shape assembly, but is limited to object-scale scenes (e.g., furniture and tableware). RAP extends Euclidean-space flow matching to large-scale, cross-domain multi-view registration and achieves superior performance by scaling the training data, performing canonicalized generation conditioned on local embeddings, and rigidifying the flow sampling at test time.
@@ 3 Generative Point Cloud Registration
Figure 2: Overview of our approach to multi-view point cloud registration. Starting with unposed point clouds $\mathcal{P}$, we sample points $\mathcal{Q}$ with corresponding local features $\mathcal{F}$. We use a diffusion transformer (DiT) with alternating-attention blocks for conditional flow matching that generates the aggregated point cloud $\hat{\mathbf{X}}(0)$ from Gaussian noise $\mathbf{X}(1)$. Finally, we recover the individual transformations $\hat{\mathbf{T}}_{i}$ using singular value decomposition (SVD) from the aggregated point cloud and apply them to the original unposed point clouds to get the registered point clouds $\hat{\mathbf{P}}^{r}$. The example illustrates submap registration on the KITTI [geiger2013ijrr] dataset.
@@ 3.1 Problem Definition
We consider the general multi-view point cloud registration problem. The input is a set of $N\geq 2$ unordered point clouds $\mathcal{P}=\{\mathbf{P}_{i}\in\mathbb{R}^{3\times M_{i}}\}_{i=1}^{N}$, where $M_{i}$ is the point count of the $i$-th point cloud. These point clouds may come from individual LiDAR or depth-camera scans, or from accumulated point cloud maps built by SLAM or photogrammetry systems. We do not assume any initial guess of the transformation from the coordinate frame of each point cloud to a global frame, but we assume that the point clouds are observations of the same scene and can be registered into a single connected point cloud, even if the overlap is small.
The goal is to estimate registered point clouds $\mathcal{P}^{r}=\{\mathbf{P}_{i}^{r}\in\mathbb{R}^{3\times M_{i}}\}_{i=1}^{N}$ in a common global coordinate frame. In practice, there may be non-rigid deformations caused by effects such as motion distortion, dynamic objects, or SLAM drift. When these non-rigid effects are negligible, the registered point clouds can be transformed from the inputs by a set of rigid body transformations $\mathcal{T}=\{\mathbf{T}_{i}\in\mathrm{SE}(3)\}_{i=1}^{N}$.
@@ 3.2 Flow Matching for Multi-view Registration
with $\Delta t=1/\kappa$. After integration, the resulting $\hat{\mathbf{X}}(0)$ approximates the registered point cloud. We then partition $\hat{\mathbf{X}}(0)$ into per-view subsets and estimate the corresponding poses $\hat{\mathbf{T}}_{i}$ via the Kabsch algorithm [arun1987pami] using SVD.
@@ 3.3 Rigidity-forcing Inference
To illustrate Euler integration with rigidity-forcing, we define the per-view projection operator $\Pi$ of an estimate $\hat{\mathbf{X}}_{i}(0)$ onto the rigid orbit of the input $\mathbf{P}_{i}$, as
@@
where $(\hat{\mathbf{R}}_{i},\hat{\mathbf{t}}_{i})$ is the optimal rigid transformation between $\mathbf{P}_{i}$ and $\hat{\mathbf{X}}_{i}(0)$, computed via the Kabsch algorithm [arun1987pami]. Given the current state $\mathbf{X}(t)$ and the velocity $\mathbf{V}_{\theta}(t,\mathbf{X}(t)\!\mid\!\mathbf{C})$, we extrapolate the registered point cloud estimate, as